Two balls A and B having masses 1 kg and 2 kg ,moving with speeds 21 m/s  and 4 m/s respectively in opposite direction ,collide head on.After collision A moves with a speed of 1 m/s in the same direction ,then the coefficient of  restitution is  ?

Grade:12

1 Answers

Rajdeep
5 years ago
HELLO THERE!
 
Let the coefficient of restitution be e.
 
We know, that
e = \frac{v_{separation}}{v_{approach}}
 
Velocity of approach:
First,we have to calculate the velocity of approach.
Since, the two bodies are moving in oppsite direction, the velocity of approach is the sum of the two velocities.
 
So, velocity of approach = 21 + 4 = 25 m/s
 
Velocity of separation:
After the collision, the question says that the block A moves in the same direction with 1 m/s, so we can get the clue that, the other block comes at rest. Hence, the velocity of separation is the velocity with which block A is moving away from B since block B is at rest!
 
Velocity of separation = 1 m/s
 
So, 
 
e = \frac{v_{separation}}{v_{approach}} \\\\=\frac{1}{25}
 
That’s your answer!